Students are traveling in two cars to a football game 135 miles away. The first car leaves on time and travels at an average speed of 45 miles per hour. The second car starts 1/2 hour later and travels at an average speed of 55 mile per hour. How long will it take the second car to catch up to the first car before the first car arrives at the game?
Found 3 solutions by htmentor, amoresroy, richwmiller:
Answer by htmentor(1343)

Students are traveling in two cars to a football game 135 miles away. The first car leaves on time and travels at an average speed of 45 miles per hour. The second car starts 1/2 hour later and travels at an average speed of 55 mile per hour. How long will it take the second car to catch up to the first car before the first car arrives at the game?
======================================================
Car 1 has a head start of 45 mi/hr * 0.5 hr = 22.5 mi
The RELATIVE speed of the two cars is 55 - 45 = 10 mi/hr
At this relative speed, the time required to make up this distance is:
22.5 mi/10 mi/hr = 2.25 hrs

Answer by amoresroy(361)

Students are traveling in two cars to a football game 135 miles away. The first car leaves on time and travels at an average speed of 45 miles per hour. The second car starts 1/2 hour later and travels at an average speed of 55 mile per hour. How long will it take the second car to catch up to the first car before the first car arrives at the game?
Let x = the number of hours it will take the second car to catch up to the first car before the first car arrives at the game
To catch-up, the second car must travel the same distance as the first car.
Where:
55x = distance traveled by the second car
45(x+.5) = distance traveled by the second car
Equation:
55x = 45(x+.5)
55x - 45x = 22.5
10x = 22.5
x = 2.25 or 2 hours and 15 minutes
It will take 2 hours and 15 minutes for the second car to catch up to the first car.

Answer by richwmiller(17219)

the first car went 22.5 miles in 30 minutes
The second car will gain on the first at 10 miles an hour
10t=22.5
t=22.5/10=2.25 hours
It will take the first car 3 hrs to go the 135 miles.
the second car, will take 55*t=135
t=135/55=2.45 hours
So they will meet
55*2.45=134.75 miles
They will meet 1/4 mile from their destination